# An Empirical Study of Mamba-based Pedestrian Attribute Recognition

### v1

@@ 4.3 Comparison on Public PAR Benchmarks
DeepMAR (ACPR 2015) [24] CaffeNet 73.79 62.02 74.92 76.21 75.56 - - - - -
HPNet (ICCV 2017) [31] Inception 76.12 65.39 77.33 78.79 78.05 - - - - -
JRL (ICCV 2017) [44] AlexNet 74.74 - 75.08 74.96 74.62 - - - - -
GRL (IJCAI 2018) [56] Inception-V3 81.20 - 77.70 80.90 79.29 - - - - -
MsVAA (ECCV 2018) [38] ResNet101 - - - - - 78.34 65.57 77.37 79.17 78.26
RA (AAAI 2019) [57] Inception-V3 81.16 - 79.45 79.23 79.34 - - - - -
VRKD (IJCAI 2019) [28] ResNet50 78.30 69.79 82.13 80.35 81.23 - - - - -
AAP (IJCAI 2019) [16] ResNet50 81.42 68.37 81.04 80.27 80.65 - - - - -
VAC (CVPR 2019) [15] ResNet50 - - - - - 79.23 64.51 75.77 79.43 77.10
ALM (ICCV 2019) [41] BN-Inception 81.87 68.17 74.71 86.48 80.16 79.79 64.79 73.93 82.03 77.77
JLAC (AAAI 2020) [40] ResNet50 83.69 69.15 79.31 82.40 80.82 79.23 64.42 75.69 79.18 77.40
SSCsoft (ICCV 2021) [18] ResNet50 82.77 68.37 75.05 87.49 80.43 - - - - -
IAA-Caps (PR 2022) [50] OSNet 81.72 68.47 79.56 82.06 80.37 - - - - -
MCFL (NCA 2022) [2] ResNet50 84.04 67.28 73.44 87.75 79.96 - - - - -
DRFormer (NC 2022) [42] ViT-B/16 81.81 70.60 80.12 82.77 81.42 - - - - -
VAC-Combine (IJCV 2022) [14] ResNet50 81.30 70.12 81.56 81.51 81.54 - - - - -
DAFL (AAAI 2022) [19] ResNet50 83.72 68.18 77.41 83.39 80.29 81.04 66.70 76.39 82.07 79.13
CGCN (TMM 2022) [6] ResNet50 84.70 54.40 60.03 83.68 70.49 - - - - -
CAS-SAL-FR (IJCV 2022) [53] ResNet50 84.18 68.59 77.56 83.81 80.56 - - - - -
VTB (TCSVT 2022) [3] ViT-B/16 82.67 69.44 78.28 84.39 80.84 81.34 67.48 76.41 83.32 79.35
PARformer (TCSVT 2023) [7] Swin-L 84.13 69.94 79.63 88.19 81.35 - - - - -
VTB* (TCSVT 2022) [3] ViT-L/14 83.69 69.78 78.09 85.21 81.10 81.36 67.58 76.19 84.00 79.52
PromptPAR (arXiv-2023) [46] ViT-L/14 85.45 71.61 79.64 86.05 82.38 83.14 69.62 77.42 85.73 81.00
SequencePAR (arXiv-2023) [21] ViT-L/14 - 71.47 82.40 82.09 82.05 - 70.14 81.37 81.22 81.10
@@
Results on WIDER dataset. As shown in Table 3, our model achieved mA of 89.38, surpassing VTB (ViT-B/16, 88.2), which further validates the feasibility of our model. Compared with the pre-trained vision-language CLIP [36] model based PAR, e.g., the VTB* and PromptPAR, the VMamba-based model is still inferior to these models. In our future works, we will consider knowledge distillation strategies to further augment the performance of VMamba-based PAR.
Table 3: Comparison with state-of-the-art methods on WIDER datasets. “V” means Vision branch.
Methods Backbone mA
R*CNN (ICCV 2015)[8] VGG16 80.5
DHC (ECCV 2016)[30] VGG16 81.3
SRN (CVPR 2017)[58] ResNet101 86.2
DIAA (ECCV 2018)[39] ResNet101 86.4
Da-HAR (AAAI 2020)[52] ResNet101 87.3
VAC-Combine (IJCV 2022) [14] ResNet50 88.4
VTB (TCSVT 2022) [3] ViT-B/16 88.2
VTB* (TCSVT 2022) [3] ViT-L/14 91.2
PromptPAR (arXiv-2023) [46] ViT-L/14 92.0
MambaPAR (V) VMamba-B 89.38
Results on PETA-ZS dataset. In addition to standard evaluations, we also evaluated the model’s performance on datasets using a zero-shot setting. According to the results reported in Table 4, our model achieved mA, Acc, Prec, Recall, and F1 metrics of 74.43, 60.56, 73.12, 74.51, and 73.37, respectively, which are close to or even better than VTB (ViT-B/16, 75.13/60.50/73.29/74.40/73.38) on certain metrics.
Results on RAP-ZS dataset. As shown in Table 4, the experimental results on the RAP-ZS dataset show that VTB based on ViT-B/16 achieved mA, Acc, Prec, Recall, and F1 metrics of 75.76, 64.73, 74.93, 80.85, and 77.35, respectively, while our experimental results were more excellent, reaching 77.34, 66.64, 75.37, 83.57, and 78.86, respectively. These results indicate that our model also performs excellently in zero-shot settings, and we believe the key reason is the high adaptability of the Mamba model to pedestrian attribute recognition.
Table 4: Comparison with state-of-the-art methods on PETA-ZS and RAP-ZS datasets. “V” means Vision branch.
@@
MsVAA (ECCV 2018) [38] ResNet101 71.53 58.67 74.65 69.42 71.94 72.04 62.13 75.67 75.81 75.74
VAC (CVPR 2019) [15] ResNet50 71.91 57.72 72.05 70.64 70.90 73.70 63.25 76.23 76.97 76.12
ALM (ICCV 2019) [41] BN-Inception 73.01 57.78 69.50 73.69 71.53 74.28 63.22 72.96 80.73 76.65
JLAC (AAAI 2020) [40] ResNet50 73.60 58.66 71.70 72.41 72.05 76.38 62.58 73.14 79.20 76.05
Jia et al. (Arxiv 2021) [20] ResNet50 71.62 58.19 73.09 70.33 71.68 72.32 63.61 76.88 76.62 76.75
MCFL (NCA 2022) [2] ResNet50 72.91 57.04 68.47 74.35 71.29 74.37 63.37 71.21 83.86 77.02
VTB (TCSVT 2022) [3] ViT-B/16 75.13 60.50 73.29 74.40 73.38 75.76 64.73 74.93 80.85 77.35
VTB* (TCSVT 2022) [3] ViT-L/14 77.18 63.12 74.77 77.24 75.50 79.17 68.34 76.81 84.51 80.07
PromptPAR (arXiv-2023) [46] ViT-L/14 80.08 66.02 76.53 80.49 77.77 80.43 70.39 78.48 85.57 81.52
SequencePAR (arXiv-2023) [21] ViT-L/14 - 66.70 78.75 78.52 78.40 - 70.28 82.13 80.55 81.14
@@
4.4 Ablation Study
@@
Effects of Vision-Semantic Fusion for PAR. As shown in Table 1, we conducted ablation experiments on the PA100K and PETA datasets for the visual semantic fusion (VSF) module. The effectiveness of this module has been validated by integrating with the Transformer networks, but not with the Mamba networks. Thus, we exploit two vision Mamba models, i.e., the VMamba-B and Vim-S, to check whether VSF still works for the Mamba-based PAR framework. It is easy to find that the pure Vim-S based model achieves (PA100K) 81.45, 74.92, 82.68, 84.27, 83.15 and (PETA) 79.28, 76.77, 85.24, 86.58, 85.44 in the five indicators of mA, Acc, Prec, Recall, and F1, respectively. After introducing the VSF module highlighted in the red dashed rectangle, the results reach 80.87, 79.33, 86.48, 88.79, 87.21 on the PA100K dataset and 84.25, 76.07, 82.73, 87.20, 84.56 on the PETA dataset, respectively, which proves the effectiveness of the semantic labels for the Vim-based PAR framework. However, we also find that similar conclusions do not hold for the VMamba-B based PAR framework, that is to say, the performance drops when adopting the semantic labels of all attributes.
@@
Transformer vs Mamba for PAR. As shown in Table 5, we compare the performance of ViT-S, Vim-S, ViT-B, and VMamba-B on the PA100K dataset using only the visual branch. One can find that the vision Mamba-based PAR framework achieves comparable or even better performance than vision Transformer based models at the same size. These comparisons demonstrate that it is still a promising research direction to adopt the Mamba for pedestrian attribute recognition.
@@
Table 5: Results of different backbones on PA100k dataset. (Only Vision branch)
@@ 4.5 Efficiency Analysis
Figure 4: Comparison on (a) model parameters, (b) Time cost, (c) GPU cost, and (d) FLOPs.
@@
Analysis of Parameters and FLOPs. As illustrated in Fig. 4(b, d), the Vim-S model exhibits a notable reduction in parameters compared to the ViT-S and ViT-B models, with no appreciable decline in performance. Additionally, the VMamba-B model displays a more compact parameter count than the ViT-B model, while maintaining comparable performance. It has been demonstrated that the Mamba-based pedestrian attribute recognition method has a reduced number of parameters and facilitates the process of fine-tuning. Furthermore, the computational costs of these methods under different batch sizes were compared. It was observed that as the batch size increased, the Transformer-based PAR method exhibited a notable reduction in computational cost. However, the computational cost of VMamba-B was significantly lower than that of ViT-B, while that of Vim-S remained largely unchanged. This demonstrates that the Mamba-based pedestrian attribute recognition method continues to exhibit reduced computation at high batch sizes and is more hardware-friendly when deployed.
4.6 Comparison of Hybrid Mamba-Transformer
In this paper, we also exploit different versions of hybrid Mamba-Transformer networks for pedestrian attribute recognition, as discussed in sub-section 3.3. We will report and discuss the experimental results of these variations in this sub-section.
Table 6: Experimental results of the proposed hybrid Mamba-Transformer networks for PAR on PETA dataset. The best and second results are highlighted in Red and Green, respectively.
@@ 4.6 Comparison of Hybrid Mamba-Transformer
(g) KDTM 81.16 79.41 87.36 87.83 87.20
(h) MaKDF 80.72 78.57 86.55 87.61 86.66
As shown in Table 6, among the six hybrid methods we proposed, method (e) achieves the best results, even outperforming the largest-scale ViT models. Others, such as (b) and (d), do not outperform ViT-B in performance, but also outperform ViT-S and Vim-S. Other than these, none of the other approaches resulted in any performance improvement, either because of design issues or because of the hybrid approach.
From the experimental results reported in Table 6, one can also find that in the method (g) KDTM, we experimented with feature distillation and predictive label distribution distillation. From the results, it seems that feature distillation has a negative effect, which may be caused by the differences between the different frameworks. Therefore, directly distilling from the predicted label distribution level can avoid the difficulty of matching the features of each layer caused by the differences between models. It can be deemed that response-level (i.e., the logits) distillation will give Vim a positive performance boost.
Based on the experimental results of KDTM, we improve the way of feature distillation and propose the method (h) MaKDF, which adds a lightweight transition module (single-layer Vim) between ViT distillation and Vim, and then adds the output of this transition module as additional information to the Vim backbone. As shown in Table 6, this assisted distillation with the Mamba module, which is independent of the backbone, brings some performance improvements, but it is still not better than the effect of direct label distribution distillation.
@@
4.7 Limitation Analysis

### v2

@@ IV-C Comparison on Public PAR Benchmarks
DeepMAR [32] ACPR15 CaffeNet 73.79 62.02 74.92 76.21 75.56 - - - - -
HPNet [15] ICCV17 Inception 76.12 65.39 77.33 78.79 78.05 - - - - -
JRL [6] ICCV17 AlexNet 74.74 - 75.08 74.96 74.62 - - - - -
GRL [7] IJCAI18 Inception-V3 81.20 - 77.70 80.90 79.29 - - - - -
MsVAA [36] ECCV18 ResNet101 - - - - - 78.34 65.57 77.37 79.17 78.26
RA [37] AAAI19 Inception-V3 81.16 - 79.45 79.23 79.34 - - - - -
VRKD [38] IJCAI19 ResNet50 78.30 69.79 82.13 80.35 81.23 - - - - -
AAP [39] IJCAI19 ResNet50 81.42 68.37 81.04 80.27 80.65 - - - - -
VAC [40] CVPR19 ResNet50 - - - - - 79.23 64.51 75.77 79.43 77.10
ALM [41] ICCV19 BN-Inception 81.87 68.17 74.71 86.48 80.16 79.79 64.79 73.93 82.03 77.77
JLAC [42] AAAI20 ResNet50 83.69 69.15 79.31 82.40 80.82 79.23 64.42 75.69 79.18 77.40
SSCNet [44] ICCV21 ResNet50 82.77 68.37 75.05 87.49 80.43 - - - - -
IAA-Caps [45] PR22 OSNet 81.72 68.47 79.56 82.06 80.37 - - - - -
MCFL [46] NCA22 ResNet50 84.04 67.28 73.44 87.75 79.96 - - - - -
DRFormer [33] NC22 ViT-B/16 81.81 70.60 80.12 82.77 81.42 - - - - -
VAC-Combine [47] IJCV22 ResNet50 81.30 70.12 81.56 81.51 81.54 - - - - -
DAFL [48] AAAI22 ResNet50 83.72 68.18 77.41 83.39 80.29 81.04 66.70 76.39 82.07 79.13
CGCN [49] TMM22 ResNet50 84.70 54.40 60.03 83.68 70.49 - - - - -
CAS-SAL-FR [50] IJCV22 ResNet50 84.18 68.59 77.56 83.81 80.56 - - - - -
VTB [9] TCSVT22 ViT-B/16 82.67 69.44 78.28 84.39 80.84 81.34 67.48 76.41 83.32 79.35
VTB* [9] TCSVT22 ViT-L/14 83.69 69.78 78.09 85.21 81.10 81.36 67.58 76.19 84.00 79.52
PARformer [8] TCSVT23 Swin-L 84.13 69.94 79.63 88.19 81.35 - - - - -
OAGCN [51] TMM23 Swin-L 87.83 69.32 78.32 87.29 82.56 - - - - -
SequencePAR [52] arXiv23 ViT-L/14 - 71.47 82.40 82.09 82.05 - 70.14 81.37 81.22 81.10
PromptPAR [34] TCSVT24 ViT-L/14 85.45 71.61 79.64 86.05 82.38 83.14 69.62 77.42 85.73 81.00
SSPNet  [53] PR24 Swin-S 83.24 70.21 80.14 82.90 81.50 - - - - -
@@
Results on PETA dataset. As shown in Table I, the evaluated MambaPAR model performs well on most evaluation metrics, with different branches and backbone networks showing varying results. Specifically, based on VMamba-B, the MambaPAR(V) achieved mA, Acc, Prec, Recall, and F1 metrics of 86.28, 80.54, 87.45, 87.95, and 87.45, respectively; the MambaPAR(V+L) achieved 85.01, 78.47, 85.12, 87.49, and 86.00. Based on Vim-S, the MambaPAR(V) achieved 81.45, 74.92, 82.68, 84.27, and 83.15; the MambaPAR(V+L) achieved 84.25, 76.07, 82.73, 87.20, and 84.56. Compared to the Transformer-based pedestrian attribute recognition models, our VMamba visual branch model surpassed VTB (ViT-B/16, 85.31/79.60/86.76/87.17/86.71) and VTB* (ViT-L/14, 86.34/79.59/86.66/87.82/86.97) in the mA, Acc, Prec, Recall, and F1 metrics. Therefore, we can conclude that our model achieves competitive results on the PETA dataset, verifying the feasibility of the Mamba model for pedestrian attribute recognition.
Results on PA100K dataset. As shown in Table I, our MambaPAR model also performs well on the PA100K dataset. Based on VMamba-B, the MambaPAR(V) achieved mA, Acc, Prec, Recall, and F1 metrics of 83.63, 81.11, 87.59, 89.9, and 88.40, respectively; the MambaPAR(V+L) achieved 81.50, 79.55, 87.54, 88.03, and 87.35. Based on Vim-S, the MambaPAR(V+L) achieved 79.28, 76.77, 85.24, 86.58, and 85.44; the MambaPAR(V) achieved 80.87, 79.33, 86.48, 88.79, and 87.21. The VMamba-B visual branch model’s performance is close to that of VTB* (85.3/81.76/87.87/90.67/88.86). These experimental results indicate that the exploited Mamba-based PAR frameworks we explored have achieved preliminary success.
Interestingly, in the experimental results of PETA and PA100K we can find that the attribute labels are useful for the Vim-S, but not for the VMamba-B. This may be because the textual representation does not match the visual representation, and the direct visual textual feature interaction using Mamba blocks at the end instead limits the representation ability of VMamba. The representation ability of Vim is weak, and the use of Mamba for feature interaction brings effective improvement.
Results on MSP60K dataset. As shown in Table II, in the random partition, our MambaPAR(V) based on VMamba-B achieved mA, Acc, Prec, Recall, and F1 metrics of 74.82, 74.68, 82.32, 86.41, and 84.18, respectively. This goes beyond SSPNet based on Swin-S. The MambaPAR(V+L) based on Vim-S achieved mA, Acc, Prec, Recall, and F1 metrics of 73.85, 73.64, 83.19, 84.29, and 83.28, respectively. The hybrid architecture(MaHDFT) based on Vim-S achieved mA, Acc, Prec, Recall, and F1 metrics of 74.08, 74.40, 82.82, 86.41, and 83.93, respectively. In the cross-domain partition, the MambaPAR(V) based on VMamba-B achieved mA, Acc, Prec, Recall, and F1 metrics of 58.26, 49.76, 62.79, 68.89, and 65.03, respectively. The MambaPAR(V+L) based on Vim-S achieved mA, Acc, Prec, Recall, and F1 metrics of 56.75, 47.34, 61.92, 64.98, and 62.80, respectively. The hybrid architecture(MaHDFT) based on Vim-S achieved mA, Acc, Prec, Recall, and F1 metrics of 58.67, 50.65, 62.39, 71.13, and 65.85, respectively. As can be seen from the results, VMamba and the hybrid architecture model has better generalization.
From the experimental results reported in the PETA, PA100K, MSP60K datasets, we can find that the VMamba-B achieves the best results over other Mamba-based PAR models. Therefore, we only report the results of pure VMamba-B based PAR framework for the rest of the datasets.
TABLE IV: Comparison with state-of-the-art methods on PETA-ZS and RAP-ZS datasets. "V" means Vision branch.
@@
MsVAA [36] ECCV18 ResNet101 71.53 58.67 74.65 69.42 71.94 72.04 62.13 75.67 75.81 75.74
VAC [40] CVPR19 ResNet50 71.91 57.72 72.05 70.64 70.90 73.70 63.25 76.23 76.97 76.12
ALM [41] ICCV19 BN-Inception 73.01 57.78 69.50 73.69 71.53 74.28 63.22 72.96 80.73 76.65
JLAC [42] AAAI20 ResNet50 73.60 58.66 71.70 72.41 72.05 76.38 62.58 73.14 79.20 76.05
Jia et al. [21] arXiv21 ResNet50 71.62 58.19 73.09 70.33 71.68 72.32 63.61 76.88 76.62 76.75
MCFL [46] NCA22 ResNet50 72.91 57.04 68.47 74.35 71.29 74.37 63.37 71.21 83.86 77.02
VTB [9] TCSVT22 ViT-B/16 75.13 60.50 73.29 74.40 73.38 75.76 64.73 74.93 80.85 77.35
VTB* [9] TCSVT22 ViT-L/14 77.18 63.12 74.77 77.24 75.50 79.17 68.34 76.81 84.51 80.07
SequencePAR [52] arXiv23 ViT-L/14 - 66.70 78.75 78.52 78.40 - 70.28 82.13 80.55 81.14
PromptPAR [34] TCSVT24 ViT-L/14 80.08 66.02 76.53 80.49 77.77 80.43 70.39 78.48 85.57 81.52
@@
Results on RAP-V1 dataset. As shown in Table III, our VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 83.12, 69.47, 78.03, 84.81, and 80.88, respectively. In comparison, the strong baseline VTB* achieved 83.69, 69.78, 78.09, 85.21, and 81.10 on these metrics, and our results are very close to theirs. Meanwhile, our model’s performance is also close to the DRFormer [33], which is also developed based on Transformer. This fully demonstrates that the performance of our Mamba-based pedestrian attribute recognition model is comparable to or even surpasses Transformer-based models.
Results on RAP-V2 dataset. As shown in Table III, our VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 81.91, 68.08, 76.41, 84.38, and 79.83, respectively, outperforming the Transformer-based VTB* (81.36/67.58/76.19/84.00/79.52). This further proves the effectiveness of our proposed pedestrian attribute recognition model.
Results on WIDER dataset. As shown in Table V, our model achieved mA of 89.38, surpassing VTB (ViT-B/16, 88.2), which further validates the feasibility of our model. Compared with the pre-trained vision-language CLIP [60] model based PAR, e.g., the VTB* and PromptPAR, the VMamba-based model is still inferior to these models. In our future works, we will consider knowledge distillation strategies to further augment the performance of VMamba-based PAR.
TABLE V: Comparison with state-of-the-art methods on WIDER datasets. “V" means Vision branch.
Methods Backbone mA
R*CNN [61] ICCV15 VGG16 80.5
DHC [li2016wider] ECCV16 VGG16 81.3
SRN [62] CVPR17 ResNet101 86.2
DIAA [36] ECCV18 ResNet101 86.4
Da-HAR [63] AAAI20 ResNet101 87.3
VAC-Combine [47] IJCV22 ResNet50 88.4
VTB [9] TCSVT22 ViT-B/16 88.2
VTB* [9] TCSVT22 ViT-L/14 91.2
PromptPAR [34] TCSVT24 ViT-L/14 92.0
MambaPAR (V) VMamba-B 89.4
Results on PETA-ZS dataset. In addition to standard evaluations, we also evaluated the model’s performance on datasets using a zero-shot setting. According to the results reported in Table IV, our model achieved mA, Acc, Prec, Recall, and F1 metrics of 74.43, 60.56, 73.12, 74.51, and 73.37, respectively, which are close to or even better than VTB (ViT-B/16, 75.13/60.50/73.29/74.40/73.38) on certain metrics.
Results on RAP-ZS dataset. As shown in Table IV, the experimental results on the RAP-ZS dataset show that VTB based on ViT-B/16 achieved mA, Acc, Prec, Recall, and F1 metrics of 75.76, 64.73, 74.93, 80.85, and 77.35, respectively, while our experimental results were more excellent, reaching 77.34, 66.64, 75.37, 83.57, and 78.86, respectively. These results indicate that our model also performs excellently in zero-shot settings, and we believe the key reason is the high adaptability of the Mamba model to pedestrian attribute recognition.
IV-D Ablation Study
@@
Effects of Vision-Semantic Fusion for PAR. As shown in Table I, we conducted ablation experiments on the PA100K and PETA datasets for the visual semantic fusion (VSF) module. The effectiveness of this module has been validated by integrating with the Transformer networks, but not with the Mamba networks. Thus, we exploit two vision Mamba models, i.e., the VMamba-B and Vim-S, to check whether VSF still works for the Mamba-based PAR framework. It is easy to find that the pure Vim-S based model achieves (PA100K) 81.45, 74.92, 82.68, 84.27, 83.15 and (PETA) 79.28, 76.77, 85.24, 86.58, 85.44 in the five indicators of mA, Acc, Prec, Recall, and F1, respectively. After introducing the VSF module highlighted in the red dashed rectangle, the results reach 80.87, 79.33, 86.48, 88.79, 87.21 on the PA100K dataset and 84.25, 76.07, 82.73, 87.20, 84.56 on the PETA dataset, respectively, which proves the effectiveness of the semantic labels for the Vim-based PAR framework. However, we also find that similar conclusions do not hold for the VMamba-B based PAR framework, that is to say, the performance drops when adopting the semantic labels of all attributes.
Figure 4: Comparison on (a) model parameters, (b) Time cost, and (c) GPU cost.
@@
Transformer vs Mamba for PAR. As shown in Table VI, we compare the performance of ViT-S, Vim-S, ViT-B, and VMamba-B on the PA100K dataset using only the visual branch. One can find that the vision Mamba-based PAR framework achieves comparable or even better performance than vision Transformer based models at the same size. These comparisons demonstrate that it is still a promising research direction to adopt the Mamba for pedestrian attribute recognition.
@@
TABLE VI: Results of different backbones on PA100k dataset. (Only Vision branch)
@@ IV-E Efficiency Analysis
Analysis of Parameters. As illustrated in Fig. 4(b), the Vim-S model exhibits a notable reduction in parameters compared to the ViT-S and ViT-B models, with no appreciable decline in performance. Additionally, the VMamba-B model displays a more compact parameter count than the ViT-B model, while maintaining comparable performance. It has been demonstrated that the Mamba-based pedestrian attribute recognition method has a reduced number of parameters and facilitates the process of fine-tuning.
IV-F Comparison of Hybrid Mamba-Transformer
In this paper, we also exploit different versions of hybrid Mamba-Transformer networks for pedestrian attribute recognition, as discussed in sub-section III-C. We will report and discuss the experimental results of these variations in this sub-section.
TABLE VII: Experimental results of the proposed hybrid Mamba-Transformer networks for PAR on PETA dataset. The best and second results are highlighted in Red and Green, respectively. "FD": Feature Distillation, "LD": Logits Distillation
@@ IV-F Comparison of Hybrid Mamba-Transformer
(g) KDTM (FD) 77.38 73.96 83.56 84.13 83.39
(g) KDTM (LD) 81.16 79.41 87.36 87.83 87.20
(h) MaKDF (FD) 80.72 78.57 86.55 87.61 86.66
@@
As presented in Table VII, our analysis of the six hybrid approaches reveals that method (e) stands out as the most effective, even surpassing the performance of the most extensive Vision Transformer (ViT) models. Additionally, methods (b) and (d), while not exceeding the capabilities of ViT-B, demonstrate superior results compared to both ViT-S and Vim-S models. Regrettably, the remaining methods failed to enhance performance, which we attribute to potential design flaws or inherent limitations of the hybrid framework itself.
From the experimental results reported in Table VII, one can also find that in the method (g) KDTM, we experimented with feature distillation and Logits distillation. From the results, it seems that feature distillation has a negative effect, which may be caused by the differences between the different frameworks. Therefore, directly distilling from the predicted label distribution level can avoid the difficulty of matching the features of each layer caused by the differences between models. It can be deemed that logits-level distillation will give Vim a positive performance boost.
Drawing insights from the KDTM experiments, we refined our feature distillation technique and introduced method (h) MaKDF. This novel approach incorporates a lightweight transition module(a single-layer Vim) situated between the ViT distillation process and the Vim model itself. The output from this intermediary module is then integrated as supplementary data into the Vim backbone. As illustrated in Table VII, this refined distillation process, facilitated by a Mamba module that operates independently of the primary model structure, achieves a measure of improvement in performance over the Baseline model. Nonetheless, it has yet to surpass the efficacy of KDTM, indicating that further refinements are necessary to maximize the potential of our hybrid models.
IV-G Visualization
As shown in the Fig. 5, most of the prediction results of the proposed MambaPAR model based on Vim-S are better, especially in the attribute recognition of knapsack. Even if the real label is not marked, it exists from the picture, and our model can predict it. The purple label in the figure obtains the attribute. This shows that our model has better generalization performance. The main errors appear in some complex scenes and the detection of pedestrian view attributes. There is great subjective uncertainty in view attributes, which may be targeted to introduce uncertainty algorithm to solve this problem.
IV-H Limitation Analysis